OpenStudy (anonymous):

please help, so confused. find the inverse for each function. f(x)=1/5+5 & x=-5

OpenStudy (anonymous):

i know for the f(x) i have to change it to y=1/5x+5 then subtract 3 from each side but not too sure how to do so.

jimthompson5910 (jim_thompson5910):

The original function f(x) is this right? \[\Large f(x) = \frac{1}{5x+5}\]

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

\[\Large f(x) = \frac{1}{5x+5}\] \[\Large y = \frac{1}{5x+5}\] \[\Large x = \frac{1}{5y+5}\] Do you see how to solve for y from here?